# AutoGUI: Scaling GUI Grounding with Automatic Functionality Annotations from LLMs

### v1

@@ Abstract
User interface understanding with vision-language models has received much attention due to its potential for enabling next-generation software automation. However, existing UI datasets either only provide large-scale context-free element annotations or contextualized functional descriptions for elements at a much smaller scale. In this work, we propose the AutoGUI pipeline for automatically annotating UI elements with detailed functionality descriptions at scale. Specifically, we leverage large language models (LLMs) to infer element functionality by comparing the UI content changes before and after simulated interactions with specific UI elements. To improve annotation quality, we propose LLM-aided rejection and verification, eliminating invalid and incorrect annotations without human labor. We construct an AutoGUI-704k dataset using the proposed pipeline, featuring multi-resolution, multi-device screenshots, diverse data domains, and detailed functionality annotations that have never been provided by previous datasets. Human evaluation shows that the AutoGUI pipeline achieves annotation correctness comparable to trained human annotators. Extensive experimental results show that our AutoGUI-704k dataset remarkably enhances VLM’s UI grounding capabilities, exhibits significant scaling effects, and outperforms existing web pre-training data types. We envision AutoGUI as a scalable pipeline for generating massive data to build GUI-oriented VLMs. AutoGUI dataset can be viewed at this anonymous URL: https://autogui-project.github.io/.
@@ 1 Introduction
User interface understanding with visual language models(VLMs) (Hong et al., 2023; Cheng et al., 2024; You et al., 2024a; Lee et al., 2023; Baechler et al., 2024) has received wide attention due to its potential in fundamentally transforming how we interact with software as well as unleashing unseen flexibility for existing apps (Fig. 1). Functionality prediction, which aims to understand the semantic purpose and interactive affordance of individual UI elements, is a crucial task that goes beyond previous UI understanding tasks focusing on structural mapping between UI code and visual layout, such as UI REG/REC (Hong et al., 2023; Li et al., 2020a) and diagram to code (Xia et al., 2024; Liu et al., 2023a).
@@
To enhance the UI understanding capability of VLMs, large-scale high-quality training data is indispensable. However, the scale of existing open-source datasets (Li et al., 2020a; Deka et al., 2017a; Li et al., 2020b; Kapoor et al., 2024; Wang et al., 2021) for UI understanding remains on the order of millions, significantly fewer than natural image datasets such as LAION-5B (Schuhmann et al., 2022). Additionally, the prevailing methods (Deka et al., 2017a; Li et al., 2020a) for collecting UI annotation are labor-intensive, leading to prohibitive costs that hinder scalability. Moreover, existing UI understanding datasets predominantly focus on describing either the visual appearance (Li et al., 2020a; b) (e.g., a button beside the navigation bar), element categories (Cheng et al., 2024) (e.g., “menu button”), or brief functions weakly related to the UI context (Bai et al., 2021) (e.g., “show more information”) shown in Fig. 2. These datasets lack contextual functional descriptions of UI elements, which poses a challenge for VLMs in comprehending the roles these elements serve within specific UI contexts, such as distinguishing between two visually similar magnifying glass icons that may represent distinct functionalities like searching and zooming.
Figure 1: UI understanding VLMs could plan a trip to the airport by integrating information across different apps and modes of transportation.
@@
To address the challenge. we propose AutoGUI, a scalable and automatic UI data annotation pipeline that provides unlimited UI element functionality annotations. Our annotation pipeline automatically collects UI interaction trajectories and leverages large language models (LLMs) to infer element functionalities based on UI content changes, eliminating the need for manual annotation by human experts. Initially, the proposed pipeline crawls a multitude of interaction trajectories on either a web browser or an Android emulator and captures screenshots at various aspect ratios. Subsequently, we use open-source LLMs (AI@Meta, 2024) to annotate the functionalities of elements on collected GUIs based on changes to UI contents when interacting with these elements. To ensure data quality, LLM-aided rejection is utilized to eliminate invalid samples, such as incompletely rendered UIs. Additionally, inspired by recent works on LLM verification (Weng et al., 2022; Lightman et al., 2023), multiple LLMs are prompted as verifiers to identify false functionality predictions. With both the rejection and verification processes, our pipeline removes unclear and invalid samples.
We curate the AutoGUI-704k dataset with the proposed pipeline. AutoGUI-704k contains 704k high-quality functionality grounding and referring tasks used to finetune and evaluate open-source VLMs. With the vast knowledge embedded within LLMs (e.g., Llama-3-70B (AI@Meta, 2024)) and fast inference infrastructure (Kwon et al., 2023; Gugger et al., 2022), our pipeline can efficiently annotate high-quality samples at a large scale and substantially reduced cost compared to traditional methods. Moreover, pioneer experiments find that our pipeline achieves annotation accuracy of 96.7% comparable to a trained human annotator.
@@
Based on the collected AutoGUI-704k dataset, we finetune open-source VLMs that own little UI grounding capabilities. Experimental results demonstrate that data collected through our AutoGUI pipeline significantly enhances the VLMs’ UI grounding accuracy and exhibits remarkable scaling effects. The results also show that our functionality annotation type is superior to the data type directly derived from web HTML code (Hong et al., 2023; Cheng et al., 2024), serving as a promising data source for building VLMs capable of UI grounding.
@@ 2 Related Works
Figure 2: Our functionality annotations vs. the annotations provided by existing UI datasets. The proposed AutoGUI (right) can generate element annotations rich in functional semantics.
@@ 2.1 Recent Advancement of VLMs
Recently, a new wave of research has started to enhance LLMs with the capability of processing both visual and textual information (Alayrac et al., 2022; Chen et al., 2023a; Li et al., 2023; Lin et al., 2023a; Liu et al., 2023b; Lin et al., 2023b; Chen et al., 2023b; Lu et al., 2024; Bai et al., 2023; Wang et al., 2024a; Zhu et al., 2024; Wang et al., 2024b; Li et al., 2024; Zhang et al., 2024a; You et al., 2024a; Laurençon et al., 2024; Peng et al., 2024; Driess et al., 2023), opening the new field of Vision Language Model (VLM). Pioneering efforts Flamingo (Alayrac et al., 2022) uses interleaved visual and language inputs as prompts and shows remarkable few-shot visual question-answering capability. Fueled by GPT-4 (Team, 2024), both academia and industry have endeavored to democratize its amazing multimodal reasoning capability. LLaVA (Liu et al., 2023b) and LLaMA-Adapter (Zhang et al., 2024a) have attempted to align vision encoders (Dosovitskiy et al., 2021) with LLMs to enable visual instruction following. Models such as VisionLLM (Wang et al., 2024b), Ferret (You et al., 2024a), and Qwen-VL (Bai et al., 2023) further enhance these capabilities with robust visual grounding. Additionally, Research is also expanding into VLM applications in scenarios rich in textual imagery (Tang et al., 2022; Ye et al., 2023b; a; Liu et al., 2024c) and embodied interactions (Driess et al., 2023; Mu et al., 2023), offering new possibilities in multimodal reasoning. Despite these advancements, the domain of UI understanding remains under-explored due to data scarcity. This paper proposes an autonomous UI annotation pipeline to tackle this challenge, aiming to expand the data available for training VLMs in this crucial area.
2.2 Existing UI Datasets and Benchmarks
Unlike mature natural image datasets (Russakovsky et al., 2014; Schuhmann et al., 2022), UI understanding datasets have received less attention in computer vision. Several efforts have been made to develop mobile UI modeling datasets (Wang et al., 2021; Li et al., 2020a; b; Bai et al., 2021; Burns et al., 2022), primarily annotating the RICO dataset (Deka et al., 2017b), which includes 72K screenshots from Android apps. Examples include Widget Captioning (Li et al., 2020a), which analyzes captions and linguistic features of UI elements, and RICOSCA (Li et al., 2020b), which maps single-step instructions to UI locations. Recently, MoTIF (Burns et al., 2022) and AITW (Rawles et al., 2023) have been proposed to focus on interpreting high-level instructions in Android environments. However, these manually curated and crowd-annotated datasets are limited in size and costly to update, presenting challenges in adapting to new UI types.
The web scenario has also gained much attention. WebShop (Yao et al., 2022), as an early attempt, introduces a simplified simulator for web navigation tasks. More recent projects, such as Mind2Web (Deng et al., 2024) and WebArena (Zhou et al., 2023), have developed realistic and reproducible web environments to improve web agent capabilities. VisualWebBench (Liu et al., 2024b) has established a comprehensive evaluation framework for VLMs, focusing on UI grounding. To tackle data insufficiency issues, recent studies like SeeClick (Cheng et al., 2024) and CogAgent (Hong et al., 2023) have utilized the latest Common Crawl data to create large-scale datasets. However, these data are derived from HTML code snippets which contain plenty of noise.
This paper aims to address the aforementioned limitations of existing UI datasets by introducing an automatic LLM-based annotation pipeline. By focusing on contextual functional descriptions of elements, our pipeline aims to enhance VLM’s capability of understanding users’ functional intents. The advantages of our AutoGUI dataset over existing datasets are summarized in Tab. 1.
3 AutoGUI: Automatic Functionality Annotation Pipeline
Table 1: Comparing our AutoGUI dataset with existing large-scale UI datasets. Multi-Res means the samples are collected on devices with various resolutions. Auto Anno. means the samples are collected autonomously. #Anno. means the number of annotated samples provided by the datasets.
@@ 3 AutoGUI: Automatic Functionality Annotation Pipeline
WebShop (Yao et al., 2022) Web 12k Web Navigation
Mind2Web (Deng et al., 2024) Web 2.4k Web Navigation
WebArena (Zhou et al., 2023) Web 812 Web Navigation
S2W (Wang et al., 2021) Mobile 112k Screen Summarization
Wid. Cap. (Li et al., 2020a) Mobile 163k Element Captioning
PixelHelp (Li et al., 2020b) Mobile 187 Element Grounding
RICOSCA (Li et al., 2020b) Mobile 295k Action Grounding
MoTIF (Burns et al., 2022) Mobile 6k Mobile Navigation
AITW (Rawles et al., 2023) Mobile 715k Mobile Navigation
RefExp (Bai et al., 2021) Mobile 20.8k Element Grounding
VWB (Liu et al., 2024b) Web 1.5k Elem. Ground & Ref.
SeeClick Web (Cheng et al., 2024) Web 271k Element Grounding
UI REC/REG (Hong et al., 2023) Web 400k Box2DOM, DOM2Box
Ferret-UI (You et al., 2024b) Mobile 250k Elem. Ground & Ref.
AutoGUI (ours) Web, Mobile 704k Functionality Ground & Ref.
@@ 3.1 Collecting UI Interaction Trajectories
Our pipeline initiates by collecting interaction trajectories, which are sequences of UI contents captured by interacting with UI elements. Each trajectory step captures all interactable elements and the accessibility tree (AXTree) that briefly outlines the UI structure, which will be used to generate functionality annotations. To amass these trajectories, we utilize the latest Common Crawl repository as the data source for web UIs and Android Emulator for mobile UIs. Note that illegal websites and Apps are excluded manually from the sources to ensure no pornographic or violent content is included in our dataset. Please refer to Sec. A.2 for collecting details and data license.
3.2 Functionality Annotation Based on UI Dynamics
Subsequently, the pipeline generates functionality annotations for elements in the collected trajectories. Interacting with an element $e$, by clicking or hovering over it, triggers content changes in the UI. In turn, these changes can be used to predict the functionality $f$ of the interacted element. For instance, if clicking an element causes new buttons to appear in a column, we can predict that the element likely functions as a dropdown menu activator (an example in Fig. D). With this observation, we utilize a capable LLM (i.e., Llama-3-70B (AI@Meta, 2024)) as a surrogate for humans to summarize an element’s functionality based on the UI content changes resulting from interaction. Concretely, we generate compact content differences for AXTrees before ($s_{t}$) and after ($s_{t+1}$) the interaction using a file-comparing library [^1] https://docs.python.org/3/library/difflib.html. Then, we prompt the LLM to thoroughly analyze the UI content changes (addition, deletion, and unchanged lines), present a detailed Chain-of-Thoughts (Wei et al., 2022) reasoning process explaining how the element affects the UI, and finally summarize the element’s functionality.
In cases where element interactions significantly transform the UI and cause lengthy differences—such as navigating to a new screen—we adjust our approach by using UI description changes instead of the AXTree differences. Specifically, we prompt the same LLM to discern the UI hierarchy, describe UI regions, and finally describe the entire UI functionality. After describing the UIs before and after the interaction, the LLM analyzes the description differences, presents reasoning, and summarizes the element’s functionality. This annotation process is formulated as:
$f=\text{LLM}(p_{\text{anno}},s_{t},s_{t+1})$ (1)
where $f$ is the predicted functionality, $p_{\text{anno}}$ is the annotation prompt (Tab. A and Tab. B). Examples of annotated elements are depicted in Fig. 4 and more annotation details are explained in Sec. A.4.
@@ 3.3 Removing Invalid Samples via LLM-Aided Rejection
The collected trajectories may contain invalid samples due to broken UIs, such as incomplete UI loading. These samples are meaningless as they contain corrupted UI content and can mislead the models trained with them.
To filter out these invalid samples, we introduce an LLM-aided rejection approach. Initially, hand-written rules are used to detect obvious broken cases, such as blank UI contents, UIs containing elements indicating content loading, and interaction targets outside of UIs. While these obvious cases constitute a large portion of the invalid samples, there are a few types that are difficult to detect with hand-written rules. For instance, interacting with a “view more” button might unexpectedly redirect the user to a login page instead of the desired information page due to website login restrictions. To identify these challenging samples, we prompt the annotating LLM to also act as a rejector. Specifically, the LLM takes the UI content changes, generated using a file-comparing library, as input, provides detailed reasoning on whether the changes are meaningful for predicting the element’s functionality, and finally outputs predictability scores ranging from 0 to 3. This process is formulated as follows:
$score=\text{LLM}(p_{\text{reject}},e,s_{t},s_{t+1})$ (2)
where $p_{\text{reject}}$ is the rejection prompt (Tab. C).
This approach ensures that clear and predictable samples receive higher scores, while those that are ambiguous or unpredictable receive lower scores. For instance, if a button labeled "Show More", upon interaction, clearly adds new content, this sample will considered to provide sufficient changes that can anticipate the content expansion functionality and will get a score of 3. Conversely, if clicking on a "View Profile" link fails to display the profile possibly due to web browser issues, this unpredictable sample will get a score less than 3.
After implementing empirical experiments, we deploy this LLM-based rejector to discard the bottom 30% of samples based on their scores to strike a balance between the elimination of invalid samples and the preservation of valid ones (More details in Sec. A.6). The samples that pass the hand-written rules and the LLM rejector are subsequently submitted for functionality annotation. Please see representative rejection examples in Fig. G.

### v2

@@ Abstract
User interface understanding with vision-language models (VLMs) has received much attention due to its potential for enhancing software automation. However, existing datasets used to build UI-VLMs either only contain large-scale context-free element annotations or contextualized functional descriptions for elements at a small scale. In this work, we propose the AutoGUI pipeline for automatically annotating UI elements with detailed functionality descriptions at scale. Specifically, we leverage large language models (LLMs) to infer element functionality by comparing UI state changes before and after simulated interactions. To improve annotation quality, we propose LLM-aided rejection and verification, eliminating invalid annotations without human labor. We construct a high-quality AutoGUI-704k dataset using the proposed pipeline, featuring diverse and detailed functionality annotations that are hardly provided by previous datasets. Human evaluation shows that we achieve annotation correctness comparable to a trained human annotator. Extensive experiments show that our dataset remarkably enhances VLM’s UI grounding capabilities and exhibits significant scaling effects. We also show the interesting potential use of our dataset in UI agent tasks. Please view our project at https://autogui-project.github.io/.
AutoGUI: Scaling GUI Grounding with Automatic Functionality Annotations from LLMs
Hongxin Li*1,2,3, Jingfan Chen*5, Jingran Su*5, Yuntao Chen4 Qing Li5, Zhaoxiang Zhang1,2,3,6 1University of Chinese Academy of Sciences (UCAS) 2New Laboratory of Pattern Recognition (NLPR), CASIA 3State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS), CASIA 4Hong Kong Institute of Science & Innovation, CASIA 5The Hong Kong Polytechnic University 6Shanghai Artificial Intelligence Laboratory
footnotetext: Equal contribution.footnotetext: Equally advising corresponding authors. E-mails: zhaoxiang.zhang@ia.ac.cn, chenyuntao08@gmail.com
@@ 1 Introduction
User interface understanding with visual language models(VLMs) (Hong et al., 2024; You et al., 2024; Wu et al., 2024) has received wide attention due to its potential in fundamentally transforming how we interact with software (Xie et al., 2024).
While recent work has made progress by employing structural mapping between UI code and visual layout, such as UI REG/REC(Hong et al., 2024; Li et al., 2020a) and layout-to-code conversion (Xia et al., 2024; Liu et al., 2023a; Baechler et al., 2024), a more critical challenge remains: understanding the semantic purpose and interactive affordance of individual UI elements, known as functionality understanding.
Figure 1: Our annotations are rich in functional semantics (bottom) compared with existing UI datasets.
@@
Accurate functionality understanding requires VLMs to possess strong element grounding capabilities - the ability to connect fine-grained visual elements with their referring expressions. To enhance this capability, large-scale training data is indispensable. However, the scale of open-source datasets with detailed element annotations (Li et al., 2020a, b; Kapoor et al., 2024; Gou et al., 2025) is unsatisfactory, significantly smaller than natural image datasets such as LAION-5B (Schuhmann et al., 2022). Additionally, traditional annotation methods (Deka et al., 2017a; Li et al., 2020a) are labor-intensive, leading to prohibitive costs that hinder scalability. Moreover, existing datasets typically focus on describing either element alt-texts (Cheng et al., 2024), or brief intents weakly related to UI context (Bai et al., 2021) shown in Fig. 1. These datasets lack contextual functional descriptions of UI elements, which poses a challenge for VLMs in comprehending the roles these elements serve within specific UI contexts, such as distinguishing between two visually similar magnifier icons that may represent distinct functionalities like searching and zooming.
@@
To address the challenge, we propose AutoGUI, an automatic annotating pipeline that provides unlimited element functionality annotations. Our pipeline collects UI interaction trajectories and leverages large language models (LLMs) to infer element functionalities based on UI state changes, eliminating the need for manual annotation by human experts. Initially, the proposed pipeline crawls a multitude of interaction trajectories on either a web browser or an Android emulator. Subsequently, we use open-source LLMs (AI@Meta, 2024) to annotate the functionalities of elements on collected GUIs based on changes to UI contents when interacting with these elements. To ensure data quality, LLM-aided rejection is utilized to eliminate invalid samples, such as incompletely rendered UIs. Additionally, inspired by LLM verification (Weng et al., 2022; Lightman et al., 2023), multiple LLMs are prompted as verifiers to identify false functionality descriptions. With both the rejection and verification processes, our pipeline removes unclear and invalid samples.
We curate the AutoGUI-704k dataset with the proposed pipeline, providing high-quality functionality grounding and referring tasks used to finetune and evaluate open-source VLMs.
Pioneer experiments find that our pipeline achieves annotation accuracy of 96.7% comparable to a trained human annotator.
@@
Based on the collected AutoGUI-704k dataset, we finetune open-source VLMs and demonstrate that our data significantly enhances the VLMs’ UI grounding accuracy and exhibits remarkable scaling effects. The results also show that our functionality annotation type is superior to the data type directly derived from web HTML code and metadata (Hong et al., 2024; Cheng et al., 2024), serving as a promising data source for building VLMs capable of UI grounding. Moreover, VLMs trained with our data can assist in GUI agent tasks by refining element grounding, which shows more potential use of our dataset.
@@ 2.1 Recent Advancement of VLMs
Recent research has enhanced LLMs with the capability of processing both visual and textual information (Alayrac et al., 2022; Liu et al., 2023b; Lin et al., 2023; Chen et al., 2024b; Lu et al., 2024; Wang et al., 2024a, b; Li et al., 2024b; Zhang et al., 2024a; You et al., 2024; Laurençon et al., 2024; Peng et al., 2024; Driess et al., 2023), opening the new field of VLM. Pioneering efforts Flamingo (Alayrac et al., 2022) uses interleaved visual and language inputs as prompts and shows few-shot visual question-answering capability.
LLaVA (Liu et al., 2023b) and LLaMA-Adapter (Zhang et al., 2024a) have attempted to align vision encoders (Dosovitskiy et al., 2021) with LLMs to enable visual instruction following. Advanced models such as InternVL (Chen et al., 2024b) and Qwen2-VL series (Wang et al., 2024a) are further equipped with impressive high-resolution and multi-lingual understanding abilities. Additionally, VLM are applied to scenarios rich in textual imagery (Ye et al., 2023b, a; Liu et al., 2024b) and embodied interactions (Driess et al., 2023; Kim et al., 2025). Despite these advancements, VLMs lag in UI understanding probably due to data scarcity. This paper contributes an automatic UI annotation pipeline to tackle this challenge, aiming to expand the data available for training VLMs in this crucial area.
Table 1: Comparing our AutoGUI dataset with existing UI grounding datasets. Multi-Res means the samples are collected on devices with various resolutions. Auto Anno. means the samples are collected autonomously. #Anno. means the number of annotated samples provided by the datasets. (Methods combining open-source individual datasets are not compared.)
@@
S2W (Wang et al., 2021) Mobile N/A 112k Screen Summarization
Wid. Cap. (Li et al., 2020a) Mobile N/A 163k Element Captioning
RICOSCA (Li et al., 2020b) Mobile N/A 295k Action Grounding
MoTIF (Burns et al., 2022) Mobile N/A 6k Mobile Navigation
RefExp (Bai et al., 2021) Mobile N/A 20.8k Element Grounding
SeeClick Web (Cheng et al., 2024) Web N/A 271k Element Grounding
MultiUI (Hong et al., 2024) Web, Mobile N/A 3M Act. & Elem. Ground
UGround-Web (Gou et al., 2025) Web Brief 1.3M Element Grounding
UI REC/REG (Hong et al., 2024) Web N/A 400k Box2DOM, DOM2Box
Ferret-UI (You et al., 2025) Mobile Brief 250k Elem. Ground & Ref.
AutoGUI (ours) Web, Mobile Contextual 704k Functionality Ground & Ref.
2.2 Existing UI Datasets
Unlike natural image datasets (Russakovsky et al., 2014; Schuhmann et al., 2022), UI understanding datasets are much smaller. Early-stage datasets (Wang et al., 2021; Li et al., 2020a, b; Bai et al., 2021; Burns et al., 2022) primarily annotate the RICO screenshot collection (Deka et al., 2017b), which includes 72K screenshots from Android apps. Examples include Widget Captioning (Li et al., 2020a), which analyzes captions and linguistic features of UI elements, and RICOSCA (Li et al., 2020b), which maps single-step instructions to UI locations. Recently, AITW (Rawles et al., 2023) and AndroidControl (Li et al., 2024a) have been proposed to focus on interpreting high-level instructions in Android environments. To increase data scale, SeeClick (Cheng et al., 2024), CogAgent (Hong et al., 2024), and OS-ATLAS (Wu et al., 2024) have utilized the UI metadata from Common Crawl webpages to produce massive element referring expressions. Several works Gou et al. (2025); Lin et al. (2024); Xu et al. (2024) also filter and combine existing datasets to produce all-in-one collections that incorporate diverse training tasks. In contrast, our AutoGUI-704k dataset contributes large-scale element functionality annotations, which convey contextual functionality semantics that are hardly provided in previous datasets. The advantages of our dataset are summarized in Tab. 1.
3 AutoGUI: Automatic Functionality Annotation Pipeline
@@ 3.1 Collecting UI Interaction Trajectories
Our pipeline initiates by collecting interaction trajectories, which are sequences of UI contents captured by interacting with UI elements. Each step captures all interactable elements and the accessibility tree (AXTree) that briefly outlines the UI structure, which will be used to annotate functionality. To amass these trajectories, we utilize the latest Common Crawl repository as the data source for web UIs and Android Emulator for mobile UIs. The open-source trajectories from AndroidControl (Li et al., 2024a) and MobileViews (Gao et al., 2024) are also included to enhance diversity. Note that illegal UIs are manually excluded from the sources. Please refer to Sec. A for collecting details and data license.
3.2 Automatic Functionality Annotation
The pipeline generates functionality annotations for elements in the collected trajectories. Interacting with an element $e$, by clicking or hovering over it, triggers UI content changes. In turn, these changes can be used to predict the functionality $f$ of the interacted element. For instance, if clicking an element causes new buttons to appear in a column, the element likely functions as a dropdown menu activator (an example in Fig. D). With this observation, we utilize a capable LLM (i.e., Llama-3-70B (AI@Meta, 2024)) as a surrogate for humans to summarize an element’s functionality based on the UI changes resulting from interaction. Concretely, we generate compact content differences for AXTrees before ($s_{t}$) and after ($s_{t+1}$) the interaction using a file-comparing library [^1] https://docs.python.org/3/library/difflib.html. Then, we prompt the LLM to analyze the UI content changes (addition, deletion, and unchanged lines), present a detailed Chain-of-Thoughts (Wei et al., 2022) reasoning process explaining how the element affects the UI, and finally summarize the element’s functionality.
In cases where element interactions significantly transform the UI and cause lengthy differences—such as navigating to a new screen—we adjust our approach by using UI description changes instead of the AXTree differences.
This annotation process is formulated as: $f=\text{LLM}(p_{\text{anno}},s_{t},s_{t+1})$ where $f$ is the predicted functionality, $p_{\text{anno}}$ is the annotation prompt (Tab. B and Tab. C). Examples are depicted in Fig. 3 and more annotation details are explained in Sec. A.5.
Figure 3: Element functionality annotations generated by the AutoGUI pipeline for both web and mobile domains.
@@ 3.3 Removing Invalid Samples via LLM-Aided Rejection
The collected trajectories may contain invalid samples due to broken UIs, such as incomplete UI loading, which can mislead the models trained with them. To filter out these invalid samples, we introduce an LLM-aided rejection approach. Initially, hand-written rules (detailed in Sec. A.6) are used to detect obvious bad cases, such as blank UIs, UIs containing elements indicating content loading, and interaction targets outside of UIs. However, a few types are difficult to detect with the rules. For instance, interacting with a “view more” button might unexpectedly redirect the user to a login page instead of the desired information page due to website login restrictions. To identify these challenging samples, we prompt the annotating LLM to also act as a rejector. Specifically, the LLM takes the UI content changes as input, provides detailed reasoning through whether the changes are meaningful for predicting the element’s functionality, and finally outputs predictability scores ranging from 0 to 3 (3 is empirically chosen for a balance between annotation efficiency and quality.). This process is formulated as follows: $score=\text{LLM}(p_{\text{reject}},e,s_{t},s_{t+1})$ where $p_{\text{reject}}$ is the rejection prompt (Tab. D).
This approach ensures that predictable samples receive higher scores, while unpredictable ones receive lower scores. For instance, if a button labeled "Show More", upon interaction, clearly adds new content, this sample will be considered to provide sufficient changes that can anticipate the content expansion functionality and will get a score of 3.
We deploy this rejector to discard the bottom 30% of samples based on score ranking to strike a balance between the elimination of invalid samples and the preservation of valid ones (Details in Sec. A.7). The samples that pass the rejection procedure are submitted for functionality annotation. Please see examples in Fig. G.